American Express Co. (AXP) pays a dividend of $1.16 a share each quarter. If you choose not to invest in American Express, you could also get a risk free interest rate of 4%. How much is a share of AXP stock worth if the dividend were expected to grow at 2% annually?
Answer #1

Given Quarterly Dividend= $1.16/share

All other information is provided in an annual basis.

So, Annual Dividend= 4*1.16= $ 4.64 (Since, there exists 4 quarters in a year)

As per the Dividend Discount Model of stock pricing, the price of a stock is given by,

P= D/(r-g), where r= rate of interest, g= growth rate of dividend, r= rate of interest, D= Dividend paid

So, P= 4.64/(.04-.02)= $ 232